OpenStudy (anonymous):

How do you change powers with negative exponents to powers with positive exponents?

OpenStudy (anonymous):

\[ x^{-a}=\frac{1}{x^a} \]

OpenStudy (anonymous):

wait what

OpenStudy (anonymous):

oh i get it now lol just change it but how did you get the 1

OpenStudy (anonymous):

\[ x^a\times x^{-a}=1 \]

OpenStudy (anonymous):

You move the number to either the numerator or denominator depending on where it originally was. The exponent becomes positive. The numbers originally on the denominator with negative exponents get moved to the numerator as positives and vice versa.